# On vector-valued functional equations with multiple recursive terms

### v2

@@ 9 On a reflected VAR(1) process
where $A$ is an $N\times N$ matrix with elements $a_{i,j}$, $i,j=1,\ldots,N$, such that $a_{i,j}\in[0,1)$, and $\tilde{Z}_{n}:=(Z_{1,n},\ldots,Z_{N,n})^{T}$. Let also $\tilde{S}_{n}:=(S_{1,n},\ldots,S_{N,n})^{T}$, $\tilde{A}_{n}:=(A_{1,n},\ldots,A_{N,n})^{T}$, and assume that $S_{j,n}$, $A_{j,n}$ are series of i.i.d. random variables, independent of anything else. From (80) is readily seen that
@@
9.1 The vector of the marginal Laplace-Stieltjes transforms
Let $F^{n}_{j}(s):=E(e^{-sF_{j,n}})$ be the LST of the distribution of a random variable $F_{j,n}$, and $\tilde{Z}^{n}(s)=(Z_{1}^{n}(s),\ldots,Z_{N}^{n}(s))^{T}$. Then,
$\begin{array}[]{rl}E(e^{-sZ_{j,n+1}})=&E(e^{-s(\sum_{k=1}^{N}a_{j,k}Z_{k,n}+S_{j,n}-A_{j,n+1})}+1-e^{-s[\sum_{k=1}^{N}a_{j,k}Z_{k,n}+S_{j,n}-A_{j,n+1}]^{-}})\vskip 5.69054pt\\ =&E(e^{-s\sum_{k=1}^{N}a_{j,k}Z_{k,n}})S_{j}(s)A_{j}(-s)+1-U_{j,n}^{-}(s),\end{array}$ (82)
@@ 9.1 The vector of the marginal Laplace-Stieltjes transforms
$U_{j,n}^{-}(s)=E(e^{-s[\sum_{k=1}^{N}a_{j,k}Z_{k,n}+S_{j,n}-A_{j,n+1}]^{-}}).$
By assuming that for each $n$, $Z_{j,n}$, $j=1,\ldots,N,$ are independent, then $Z^{n}(s_{1},\ldots,s_{N}):=E(e^{-\sum_{k=1}^{N}Z_{k,n}s_{k}})=\prod_{k=1}^{N}Z^{n}_{k}(s_{k})$. Now, by letting $n\to\infty$, (82) is rewritten for $j=1,\ldots,N$, $Re(s)=0$, as
@@
with $\circ$” denotes the Hadamard product, $P_{i,j}$ be an $N\times N$ matrix with the $(i,j)$ element is equal to 1, and all others equal to 0, and $K(s):=diag(K_{1}(s),\ldots,K_{N}(s))$, $L(s):=diag(L_{1}(s),\ldots,L_{N}(s))$. Note that (88), although seems to has a similar form as the one in (34), it is far more complicated since in any iterating step we have to appropriately substitute for any $\tilde{Z}(a_{i,j}s)$, $i,j=1,\ldots,N$, thus rapidly increasing the number of terms.
@@
For $s=0$, we have that $c_{0,j}=0$, $j=1,\ldots,N$. Thus, for $Re(s)\geq 0$,
$Z_{j}(s)=Z_{j}(a_{j}s)S_{j}(s)\frac{\lambda_{j}}{\lambda_{j}-s}-\frac{sc_{1,j}}{\lambda_{j}-s}$ (92)
or equivalently, in matrix notation
$\tilde{Z}(s)=[S(s)\Lambda\circ I]\sum_{k=1}^{N}\tilde{P}^{(k)}\tilde{Z}(a_{k}s)+\tilde{V}(s),$ (93)
where $\circ"$ denotes the Hadamard product, $S(s)=diag(S_{1}(s),\ldots,S_{N}(s))$, $\Lambda=diag(\lambda_{1},\ldots,\lambda_{N})$, $\tilde{V}(s):=(I-A(-s))\tilde{c}$, $A(s):=diag(A_{1}(s),\ldots,A_{N}(s))$, $\tilde{c}=(c_{1,1},\ldots,c_{1,N})^{T}$. Note that (93) has the same form as (34), where $H(s)$ is equal now to $S(s)\Lambda\circ I$.
9.2 The joint Laplace-Stieltjes transform
Note that in order to come up with the expression in (83), we assumed that for each $n$, $Z_{j,n}$, $j=1,\ldots,N$ are independent. That assumption, although it creates a somehow more complicated vector valued functional equation in (88) (see also (93)), it can be treated similarly as those in the previous sections. In the following, we consider the joint LST of the stationary vector $(Z_{1},Z_{2})^{T}$, i.e., $f(s_{1},s_{2}):=E(e^{-s_{1}Z_{1}-s_{2}Z_{2}})$, and further drop the assumption of independence, but to keep the paper readable, we assume that $N=2$. Then,
$\begin{array}[]{rl}f(s_{1},s_{2}):=&E(e^{-s_{1}Z_{1}-s_{2}Z_{2}})=E(e^{-s_{1}[\sum_{j=1}^{2}a_{1,j}Z_{j}+S_{1}-A_{1}]^{+}-s_{2}[\sum_{j=1}^{2}a_{2,j}Z_{j}+S_{2}-A_{2}]^{+}})\vskip 5.69054pt\\ =&E(\left(e^{-s_{1}[\sum_{j=1}^{2}a_{1,j}Z_{j}+S_{1}-A_{1}]}+1-e^{-s_{1}[\sum_{j=1}^{2}a_{1,j}Z_{j}+S_{1}-A_{1}]^{-}}\right)\vskip 5.69054pt\\ &\times\left(e^{-s_{2}[\sum_{j=1}^{2}a_{2,j}Z_{j}+S_{2}-A_{2}]}+1-e^{-s_{2}[\sum_{j=1}^{2}a_{2,j}Z_{j}+S_{2}-A_{2}]^{+}}\right))\vskip 5.69054pt\\ =&E(e^{-s_{1}[\sum_{j=1}^{2}a_{1,j}Z_{j}+S_{1}-A_{1}]-s_{2}[\sum_{j=1}^{2}a_{2,j}Z_{j}+S_{2}-A_{2}]})+E(e^{-s_{1}[\sum_{j=1}^{2}a_{1,j}Z_{j}+S_{1}-A_{1}]})\vskip 5.69054pt\\ &-E(e^{-s_{1}[\sum_{j=1}^{2}a_{1,j}Z_{j}+S_{1}-A_{1}]-s_{2}[\sum_{j=1}^{2}a_{2,j}Z_{j}+S_{2}-A_{2}]^{-}})+E(e^{-s_{2}[\sum_{j=1}^{2}a_{2,j}Z_{j}+S_{2}-A_{2}]})\vskip 5.69054pt\\ &+1-E(e^{-s_{2}[\sum_{j=1}^{2}a_{2,j}Z_{j}+S_{2}-A_{2}]^{-}})-E(e^{-s_{1}[\sum_{j=1}^{2}a_{1,j}Z_{j}+S_{1}-A_{1}]^{-}-s_{2}[\sum_{j=1}^{2}a_{2,j}Z_{j}+S_{2}-A_{2}]})\vskip 5.69054pt\\ &-E(e^{-s_{1}[\sum_{j=1}^{2}a_{1,j}Z_{j}+S_{1}-A_{1}]^{-}})+E(e^{-s_{1}[\sum_{j=1}^{2}a_{1,j}Z_{j}+S_{1}-A_{1}]^{-}-s_{2}[\sum_{j=1}^{2}a_{2,j}Z_{j}+S_{2}-A_{2}]^{-}}).\end{array}$ (94)
@@ 9.2 The joint Laplace-Stieltjes transform
$\begin{array}[]{rl}E(e^{-s_{1}[\sum_{j=1}^{2}a_{1,j}Z_{j}+S_{1}-A_{1}]-s_{2}[\sum_{j=1}^{2}a_{2,j}Z_{j}+S_{2}-A_{2}]})=&E(e^{-Z_{1}\sum_{j=1}^{2}a_{j,1}s_{j}-Z_{2}\sum_{j=1}^{2}a_{j,2}s_{j}})\frac{\lambda_{1}\lambda_{2}S_{1}(s_{1})S_{2}(s_{2})}{(\lambda_{1}-s_{1})(\lambda_{2}-s_{2})}\vskip 5.69054pt\\ =&\frac{\lambda_{1}\lambda_{2}S_{1}(s_{1})S_{2}(s_{2})}{(\lambda_{1}-s_{1})(\lambda_{2}-s_{2})}f(\sum_{j=1}^{2}a_{j,1}s_{j},\sum_{j=1}^{2}a_{j,2}s_{j}),\vskip 5.69054pt\\ E(e^{-s_{k}[\sum_{j=1}^{2}a_{k,j}Z_{j}+S_{k}-A_{k}]})=&\frac{\lambda_{k}S_{k}(s_{k})}{\lambda_{k}-s}f(a_{k,1}s_{k},a_{k,2}s_{k}),\,k=1,2.\end{array}$
@@
$E(e^{-s_{k}[\sum_{j=1}^{2}a_{k,j}Z_{j}+S_{k}-A_{k}]^{-}})=1+\frac{s_{k}S_{k}(\lambda_{k})}{\lambda_{k}-s_{k}}f(\lambda_{k}a_{k,1},\lambda_{k}a_{k,2}),\,k=1,2,$
@@
$\begin{array}[]{r}E(e^{-s_{1}[\sum_{j=1}^{2}a_{1,j}Z_{j}+S_{1}-A_{1}]-s_{2}[\sum_{j=1}^{2}a_{2,j}Z_{j}+S_{2}-A_{2}]^{-}})=E(e^{-s_{1}[\sum_{j=1}^{2}a_{1,j}Z_{j}+S_{1}-A_{1}]}(1+\frac{s_{2}}{\lambda_{2}-s_{2}}e^{-\lambda_{2}[\sum_{j=1}^{2}a_{2,j}Z_{j}+S_{2}]}))\vskip 5.69054pt\\ =E(e^{-s_{1}[\sum_{j=1}^{2}a_{1,j}Z_{j}+S_{1}-A_{1}]})+\frac{s_{2}}{\lambda_{2}-s_{2}}E(e^{-s_{1}[\sum_{j=1}^{2}a_{1,j}Z_{j}+S_{1}-A_{1}]-\lambda_{2}[\sum_{j=1}^{2}a_{2,j}Z_{j}+S_{2}]})\vskip 5.69054pt\\ =\frac{\lambda_{1}S_{1}(s_{1})}{\lambda_{1}-s_{1}}f(a_{1,1}s_{1},a_{1,2}s_{1})+\frac{\lambda_{1}s_{2}S_{2}(\lambda_{2})S_{1}(s_{1})}{(\lambda_{1}-s_{1})(\lambda_{2}-s_{2})}f(s_{1}a_{1,1}+\lambda_{2}a_{2,1},s_{1}a_{1,2}+\lambda_{2}a_{2,2}),\end{array}$
@@
$\begin{array}[]{c}E(e^{-s_{1}[\sum_{j=1}^{2}a_{1,j}Z_{j}+S_{1}-A_{1}]^{-}-s_{2}[\sum_{j=1}^{2}a_{2,j}Z_{j}+S_{2}-A_{2}]})\vskip 5.69054pt\\ =\frac{\lambda_{2}S_{2}(s_{2})}{\lambda_{2}-s_{2}}f(a_{2,1}s_{2},a_{2,2}s_{2})+\frac{\lambda_{2}s_{1}S_{1}(\lambda_{1})S_{2}(s_{2})}{(\lambda_{1}-s_{1})(\lambda_{2}-s_{2})}f(\lambda_{1}a_{1,1}+s_{2}a_{2,1},\lambda_{1}a_{1,2}+s_{2}a_{2,2}).\end{array}$
@@
$\begin{array}[]{l}E(e^{-s_{1}[\sum_{j=1}^{2}a_{1,j}Z_{j}+S_{1}-A_{1}]^{-}-s_{2}[\sum_{j=1}^{2}a_{2,j}Z_{j}+S_{2}-A_{2}]^{-}})\vskip 5.69054pt\\ =E(\left(1+\frac{s_{1}}{\lambda_{1}-s_{1}}e^{-\lambda_{1}[\sum_{j=1}^{2}a_{1,j}Z_{j}+S_{1}]}\right)\left(1+\frac{s_{2}}{\lambda_{2}-s_{2}}e^{-\lambda_{2}[\sum_{j=1}^{2}a_{2,j}Z_{j}+S_{2}]}\right))\vskip 5.69054pt\\ =1+\frac{s_{2}S_{2}(\lambda_{2})}{\lambda_{2}-s_{2}}f(\lambda_{2}a_{2,1},\lambda_{2}a_{2,2})+\frac{s_{1}S_{1}(\lambda_{1})}{\lambda_{1}-s_{1}}f(\lambda_{1}a_{1,1},\lambda_{1}a_{1,2})+\frac{s_{1}s_{2}S_{1}(\lambda_{1})S_{2}(\lambda_{2})}{(\lambda_{1}-s_{1})(\lambda_{2}-s_{2})}f(\lambda_{1}a_{1,1}+\lambda_{2}a_{2,1},\lambda_{1}a_{1,2}+\lambda_{2}a_{2,2}).\end{array}$
@@
$\begin{array}[]{rl}f(s_{1},s_{2})=&c_{1}(s_{1},s_{2})f(\sum_{j=1}^{2}a_{j,1}s_{j},\sum_{j=1}^{2}a_{j,2}s_{j})-c_{2}(s_{1},s_{2})f(s_{1}a_{1,1}+\lambda_{2}a_{2,1},s_{1}a_{1,2}+\lambda_{2}a_{2,2})\vskip 5.69054pt\\ &-c_{3}(s_{1},s_{2})f(\lambda_{1}a_{1,1}+s_{2}a_{2,1},\lambda_{1}a_{1,2}+s_{2}a_{2,2})+c_{4}(s_{1},s_{2})f(\sum_{j=1}^{2}a_{j,1}\lambda_{j},\sum_{j=1}^{2}a_{j,2}\lambda_{j}),\end{array}$ (95)
where
$\begin{array}[]{rl}c_{1}(s_{1},s_{2})=\frac{\lambda_{1}\lambda_{2}S_{1}(s_{1})S_{2}(s_{2})}{(\lambda_{1}-s_{1})(\lambda_{2}-s_{2})},&c_{2}(s_{1},s_{2})=\frac{\lambda_{1}s_{2}S_{1}(s_{1})S_{2}(\lambda_{2})}{(\lambda_{1}-s_{1})(\lambda_{2}-s_{2})},\vskip 5.69054pt\\ c_{3}(s_{1},s_{2})=\frac{s_{1}\lambda_{2}S_{1}(\lambda_{1})S_{2}(s_{2})}{(\lambda_{1}-s_{1})(\lambda_{2}-s_{2})},&c_{4}(s_{1},s_{2})=\frac{s_{1}s_{2}S_{1}(\lambda_{1})S_{2}(\lambda_{2})}{(\lambda_{1}-s_{1})(\lambda_{2}-s_{2})}.\end{array}$
[…]
@@
where from hereon we denote $A_{i,j}:=f(a_{1}^{i}\lambda_{1},a_{2}^{j}\lambda_{2})$, $i,j=1,2,\ldots$. Then, iterating (96) n1n-1 times results in

### v3

@@ 4.2 On a reflected VAR(1) process
where $A$ is an $N\times N$ matrix with elements $a_{i,j}$, $i,j=1,\ldots,N$, such that $a_{i,j}\in[0,1)$, and $\tilde{Z}_{n}:=(Z_{1,n},\ldots,Z_{N,n})^{T}$. Let also $\tilde{S}_{n}:=(S_{1,n},\ldots,S_{N,n})^{T}$, $\tilde{A}_{n}:=(A_{1,n},\ldots,A_{N,n})^{T}$, and assume that $S_{j,n}$, $A_{j,n}$ are series of i.i.d. random variables, independent of anything else. From (87) is readily seen that (element-wise form)
@@
4.2.1 The vector of the marginal Laplace-Stieltjes transforms
Let $F^{n}_{j}(s):=\mathbb{E}(e^{-sF_{j,n}})$ be the LST of the distribution of a random variable $F_{j,n}$, and $\tilde{Z}^{n}(s)=(z_{1}^{n}(s),\ldots,z_{N}^{n}(s))^{T}$. Then,
$\begin{array}[]{rl}\mathbb{E}(e^{-sZ_{j,n+1}})=&\mathbb{E}(e^{-s(\sum_{k=1}^{N}a_{j,k}Z_{k,n}+S_{j,n}-A_{j,n+1})}+1-e^{-s[\sum_{k=1}^{N}a_{j,k}Z_{k,n}+S_{j,n}-A_{j,n+1}]^{-}})\vskip 5.69054pt\\ =&\mathbb{E}(e^{-s\sum_{k=1}^{N}a_{j,k}Z_{k,n}})S_{j}(s)A_{j}(-s)+1-u_{j,n}^{-}(s),\end{array}$ (89)
@@ 4.2.1 The vector of the marginal Laplace-Stieltjes transforms
$u_{j,n}^{-}(s)=\mathbb{E}(e^{-s[\sum_{k=1}^{N}a_{j,k}Z_{k,n}+S_{j,n}-A_{j,n+1}]^{-}}).$
By assuming that for each $n$, $Z_{j,n}$, $j=1,\ldots,N,$ are independent, then $z^{n}(s_{1},\ldots,s_{N}):=\mathbb{E}(e^{-\sum_{k=1}^{N}Z_{k,n}s_{k}})=\prod_{k=1}^{N}z^{n}_{k}(s_{k})$. Now, by letting $n\to\infty$, (89) is rewritten for $j=1,\ldots,N$, $Re(s)=0$, as
@@
with $\circ$” denotes the Hadamard product, $P_{i,j}$ be an $N\times N$ matrix with the $(i,j)$ element is equal to 1, and all others equal to 0, and $K(s):=diag(k_{1}(s),\ldots,k_{N}(s))$, $L(s):=diag(l_{1}(s),\ldots,l_{N}(s))$. Note that (95), although seems to has a similar form as the one in (1), it is more complicated since in any iterating step we have to appropriately substitute for any $\tilde{Z}(a_{i,j}s)$, $i,j=1,\ldots,N$, thus rapidly increasing the number of terms.
Note that although (95) involves Hadamard products and multiple indices, it retains the key structural properties that are required in Theorem 1. More precisely, the arguments contract to zero, and the multiplicative term satisfies $K(0)=I$, and $\tilde{L}(0)=\tilde{0}$. Moreover, since the Hadamard product acts componentwise, it does not introduce any interaction between different coordinates. In particular, for any matrix norm (e.g., $\|A\|_{\max}:=\max_{i,j}\{|A_{i,j}|\}$), we have $\|A\circ B\|\leq\|A\|\,\|B\|$, so that the Hadamard product preserves boundedness of the iterates.
Furthermore, after $k$ recursions, the argument should be $s_{k}=a_{i_{1},j_{1}}\ldots a_{i_{k},j_{k}}s$ ($s_{0}=s$), and therefore $|s_{k}|\leq(a^{*})^{k}|s|$ for $a^{*}:=\max_{i,j}\{a_{i,j}\}<1$. Since $l_{j}(s)=O(s)$ near 0, each occurrence of $\tilde{L}(s_{k})$ introduces a factor of order $(a^{*})^{k}$, while the multiplicative factors involving $K(s_{k})$ remain uniformly bounded. Consequently, each term in the expansion is bounded by a geometric sequence of order $(a^{*})^{k}$, and the Hadamard product does not affect this decay. Consequently, the associated series converges absolutely, and the solution can be obtained by an analogous iterative procedure. So, Theorem 1 can be adapted to that case.
@@
For $s=0$, we have $c_{0,j}=0$, $j=1,\ldots,N$. For $s=\lambda_{j}$, we have $c_{1,j}=-z_{j}(a_{j}\lambda_{j})S_{j}(\lambda_{j})\lambda_{j}$. Thus, for $Re(s)\geq 0$,
$z_{j}(s)=z_{j}(a_{j}s)S_{j}(s)\frac{\lambda_{j}}{\lambda_{j}-s}-\frac{sc_{1,j}}{\lambda_{j}-s}.$ (99)
Note that the diagonal form of $A$ reduces the analysis for the vector of marginal LSTs to the scalar version investigated in [12]. In particular, the solution of (99) is
$z_{j}(s)=c_{1,j}\sum_{m=0}^{\infty}\frac{a_{j}^{m}s}{\lambda_{j}-a_{j}^{m}s}\prod_{k=0}^{m-1}\frac{\lambda_{j}S_{j}^{*}(a_{j}^{k}s)}{\lambda_{j}-a_{j}^{k}s}+\prod_{m=0}^{\infty}\frac{\lambda_{j}S_{j}^{*}(a_{j}^{m}s)}{\lambda_{j}-a_{j}^{m}s}.$ (100)
For $s=a_{j}\lambda_{j}$ in (100) we find an expression for $z_{j}(a_{j}\lambda_{j})$, and substituting in $c_{1,j}=-z_{j}(a_{j}\lambda_{j})S_{j}(\lambda_{j})\lambda_{j}$, all $c_{1,j}$, $j=1,2,\ldots,N$ are known. Then, $z_{j}(s)$ are given explicitly in (100).
In matrix notation, (99) is written as
$\tilde{Z}(s)=S(s)A(-s)\sum_{k=1}^{N}\tilde{P}^{(k)}\tilde{Z}(a_{k}s)+\tilde{V}(s),$ (101)
where $S(s)=diag(S_{1}(s),\ldots,S_{N}(s))$, $\Lambda=diag(\lambda_{1},\ldots,\lambda_{N})$, $\tilde{V}(s):=(I-A(-s))\tilde{c}$, $A(s):=diag(A_{1}(s),\ldots,A_{N}(s))$, $\tilde{c}=(c_{1,1},\ldots,c_{1,N})^{T}$. Note that (101) has the same form as (20), where $H(s)$ is equal now to S(s)A(s)S(s)A(-s), however, its diagonal form considerably simplifies the analysis (note that $H(0)=I$, $\tilde{V}(0)=\tilde{0}$).
4.2.2 The joint Laplace-Stieltjes transform
Note that in order to come up with the expression in (90), we assumed that for each $n$, $Z_{j,n}$, $j=1,\ldots,N$ are independent. That assumption, although it creates a somehow more complicated vector-valued functional equation in (95) (see also (101)), can be treated similarly to those in the previous sections. In the following, we consider the joint LST of the stationary vector, drop the assumption of independence, and to enhance readability, we focus on the case $N=2$. Thus, we aim to obtain the joint LST of the stationary vector $(Z_{1},Z_{2})^{T}$, i.e., $f(s_{1},s_{2}):=\mathbb{E}(e^{-s_{1}Z_{1}-s_{2}Z_{2}})$. Then,
$\begin{array}[]{rl}f(s_{1},s_{2}):=&\mathbb{E}(e^{-s_{1}Z_{1}-s_{2}Z_{2}})=E(e^{-s_{1}[\sum_{j=1}^{2}a_{1,j}Z_{j}+S_{1}-A_{1}]^{+}-s_{2}[\sum_{j=1}^{2}a_{2,j}Z_{j}+S_{2}-A_{2}]^{+}})\vskip 5.69054pt\\ =&\mathbb{E}(\left(e^{-s_{1}[\sum_{j=1}^{2}a_{1,j}Z_{j}+S_{1}-A_{1}]}+1-e^{-s_{1}[\sum_{j=1}^{2}a_{1,j}Z_{j}+S_{1}-A_{1}]^{-}}\right)\vskip 5.69054pt\\ &\times\left(e^{-s_{2}[\sum_{j=1}^{2}a_{2,j}Z_{j}+S_{2}-A_{2}]}+1-e^{-s_{2}[\sum_{j=1}^{2}a_{2,j}Z_{j}+S_{2}-A_{2}]^{+}}\right))\vskip 5.69054pt\\ =&\mathbb{E}(e^{-s_{1}[\sum_{j=1}^{2}a_{1,j}Z_{j}+S_{1}-A_{1}]-s_{2}[\sum_{j=1}^{2}a_{2,j}Z_{j}+S_{2}-A_{2}]})+\mathbb{E}(e^{-s_{1}[\sum_{j=1}^{2}a_{1,j}Z_{j}+S_{1}-A_{1}]})\vskip 5.69054pt\\ &-\mathbb{E}(e^{-s_{1}[\sum_{j=1}^{2}a_{1,j}Z_{j}+S_{1}-A_{1}]-s_{2}[\sum_{j=1}^{2}a_{2,j}Z_{j}+S_{2}-A_{2}]^{-}})+\mathbb{E}(e^{-s_{2}[\sum_{j=1}^{2}a_{2,j}Z_{j}+S_{2}-A_{2}]})\vskip 5.69054pt\\ &+1-\mathbb{E}(e^{-s_{2}[\sum_{j=1}^{2}a_{2,j}Z_{j}+S_{2}-A_{2}]^{-}})-\mathbb{E}(e^{-s_{1}[\sum_{j=1}^{2}a_{1,j}Z_{j}+S_{1}-A_{1}]^{-}-s_{2}[\sum_{j=1}^{2}a_{2,j}Z_{j}+S_{2}-A_{2}]})\vskip 5.69054pt\\ &-\mathbb{E}(e^{-s_{1}[\sum_{j=1}^{2}a_{1,j}Z_{j}+S_{1}-A_{1}]^{-}})+\mathbb{E}(e^{-s_{1}[\sum_{j=1}^{2}a_{1,j}Z_{j}+S_{1}-A_{1}]^{-}-s_{2}[\sum_{j=1}^{2}a_{2,j}Z_{j}+S_{2}-A_{2}]^{-}}).\end{array}$ (102)
@@ 4.2.2 The joint Laplace-Stieltjes transform
$\begin{array}[]{rl}\mathbb{E}(e^{-s_{1}[\sum_{j=1}^{2}a_{1,j}Z_{j}+S_{1}-A_{1}]-s_{2}[\sum_{j=1}^{2}a_{2,j}Z_{j}+S_{2}-A_{2}]})=&\mathbb{E}(e^{-Z_{1}\sum_{j=1}^{2}a_{j,1}s_{j}-Z_{2}\sum_{j=1}^{2}a_{j,2}s_{j}})\frac{\lambda_{1}\lambda_{2}S_{1}(s_{1})S_{2}(s_{2})}{(\lambda_{1}-s_{1})(\lambda_{2}-s_{2})}\vskip 5.69054pt\\ =&\frac{\lambda_{1}\lambda_{2}S_{1}(s_{1})S_{2}(s_{2})}{(\lambda_{1}-s_{1})(\lambda_{2}-s_{2})}f(a_{1,1}s_{1}+a_{2,1}s_{2},a_{1,2}s_{1}+a_{2,2}s_{2}),\vskip 5.69054pt\\ \mathbb{E}(e^{-s_{k}[\sum_{j=1}^{2}a_{k,j}Z_{j}+S_{k}-A_{k}]})=&\frac{\lambda_{k}S_{k}(s_{k})}{\lambda_{k}-s}f(a_{k,1}s_{k},a_{k,2}s_{k}),\,k=1,2.\end{array}$
@@
$\mathbb{E}(e^{-s_{k}[\sum_{j=1}^{2}a_{k,j}Z_{j}+S_{k}-A_{k}]^{-}})=1+\frac{s_{k}S_{k}(\lambda_{k})}{\lambda_{k}-s_{k}}f(\lambda_{k}a_{k,1},\lambda_{k}a_{k,2}),\,k=1,2,$
@@
$\begin{array}[]{r}\mathbb{E}(e^{-s_{1}[\sum_{j=1}^{2}a_{1,j}Z_{j}+S_{1}-A_{1}]-s_{2}[\sum_{j=1}^{2}a_{2,j}Z_{j}+S_{2}-A_{2}]^{-}})=\mathbb{E}(e^{-s_{1}[\sum_{j=1}^{2}a_{1,j}Z_{j}+S_{1}-A_{1}]}(1+\frac{s_{2}}{\lambda_{2}-s_{2}}e^{-\lambda_{2}[\sum_{j=1}^{2}a_{2,j}Z_{j}+S_{2}]}))\vskip 5.69054pt\\ =\mathbb{E}(e^{-s_{1}[\sum_{j=1}^{2}a_{1,j}Z_{j}+S_{1}-A_{1}]})+\frac{s_{2}}{\lambda_{2}-s_{2}}E(e^{-s_{1}[\sum_{j=1}^{2}a_{1,j}Z_{j}+S_{1}-A_{1}]-\lambda_{2}[\sum_{j=1}^{2}a_{2,j}Z_{j}+S_{2}]})\vskip 5.69054pt\\ =\frac{\lambda_{1}S_{1}(s_{1})}{\lambda_{1}-s_{1}}f(a_{1,1}s_{1},a_{1,2}s_{1})+\frac{\lambda_{1}s_{2}S_{2}(\lambda_{2})S_{1}(s_{1})}{(\lambda_{1}-s_{1})(\lambda_{2}-s_{2})}f(s_{1}a_{1,1}+\lambda_{2}a_{2,1},s_{1}a_{1,2}+\lambda_{2}a_{2,2}),\end{array}$
@@
$\begin{array}[]{l}\mathbb{E}(e^{-s_{1}[\sum_{j=1}^{2}a_{1,j}Z_{j}+S_{1}-A_{1}]^{-}-s_{2}[\sum_{j=1}^{2}a_{2,j}Z_{j}+S_{2}-A_{2}]})\vskip 5.69054pt\\ =\frac{\lambda_{2}S_{2}(s_{2})}{\lambda_{2}-s_{2}}f(a_{2,1}s_{2},a_{2,2}s_{2})+\frac{\lambda_{2}s_{1}S_{1}(\lambda_{1})S_{2}(s_{2})}{(\lambda_{1}-s_{1})(\lambda_{2}-s_{2})}f(\lambda_{1}a_{1,1}+s_{2}a_{2,1},\lambda_{1}a_{1,2}+s_{2}a_{2,2}).\end{array}$
@@
$\begin{array}[]{l}\mathbb{E}(e^{-s_{1}[\sum_{j=1}^{2}a_{1,j}Z_{j}+S_{1}-A_{1}]^{-}-s_{2}[\sum_{j=1}^{2}a_{2,j}Z_{j}+S_{2}-A_{2}]^{-}})\vskip 5.69054pt\\ =\mathbb{E}(\left(1+\frac{s_{1}}{\lambda_{1}-s_{1}}e^{-\lambda_{1}[\sum_{j=1}^{2}a_{1,j}Z_{j}+S_{1}]}\right)\left(1+\frac{s_{2}}{\lambda_{2}-s_{2}}e^{-\lambda_{2}[\sum_{j=1}^{2}a_{2,j}Z_{j}+S_{2}]}\right))\vskip 5.69054pt\\ =1+\frac{s_{2}S_{2}(\lambda_{2})}{\lambda_{2}-s_{2}}f(\lambda_{2}a_{2,1},\lambda_{2}a_{2,2})+\frac{s_{1}S_{1}(\lambda_{1})}{\lambda_{1}-s_{1}}f(\lambda_{1}a_{1,1},\lambda_{1}a_{1,2})+\frac{s_{1}s_{2}S_{1}(\lambda_{1})S_{2}(\lambda_{2})}{(\lambda_{1}-s_{1})(\lambda_{2}-s_{2})}f(\lambda_{1}a_{1,1}+\lambda_{2}a_{2,1},\lambda_{1}a_{1,2}+\lambda_{2}a_{2,2}).\end{array}$
@@
$\begin{array}[]{rl}f(s_{1},s_{2})=&c_{1}(s_{1},s_{2})f(T(s_{1},s_{2}))-c_{2}(s_{1},s_{2})f(T(s_{1},\lambda_{2}))\vskip 5.69054pt\\ &-c_{3}(s_{1},s_{2})f(T(\lambda_{1},s_{2}))+c_{4}(s_{1},s_{2})f(T(\lambda_{1},\lambda_{2})),\end{array}$ (103)
where $T(s_{1},s_{2}):=A^{T}\tilde{s}=(\sum_{j=1}^{2}a_{j,1}s_{j},\sum_{j=1}^{2}a_{j,2}s_{j})$, $\tilde{s}:=(s_{1},s_{2})^{T}$, and
$\begin{array}[]{rl}c_{1}(s_{1},s_{2})=\frac{\lambda_{1}\lambda_{2}S_{1}(s_{1})S_{2}(s_{2})}{(\lambda_{1}-s_{1})(\lambda_{2}-s_{2})},&c_{2}(s_{1},s_{2})=\frac{\lambda_{1}s_{2}S_{1}(s_{1})S_{2}(\lambda_{2})}{(\lambda_{1}-s_{1})(\lambda_{2}-s_{2})},\vskip 5.69054pt\\ c_{3}(s_{1},s_{2})=\frac{s_{1}\lambda_{2}S_{1}(\lambda_{1})S_{2}(s_{2})}{(\lambda_{1}-s_{1})(\lambda_{2}-s_{2})},&c_{4}(s_{1},s_{2})=\frac{s_{1}s_{2}S_{1}(\lambda_{1})S_{2}(\lambda_{2})}{(\lambda_{1}-s_{1})(\lambda_{2}-s_{2})}.\end{array}$ (104)
Our aim is to solve the functional equation (103) iteratively. A crucial step is to verify that the mapping $T(s_{1},s_{2}):=A^{T}\tilde{s}$ is a contraction. Although we assume that $a_{i,j}\in[0,1)$, this alone is not sufficient to guarantee convergence. In order to ensure that the iterates $T^{(n)}(s_{1},s_{2})$ converge to the unique fixed point $(0,0)$, it is required that $T$ is a contraction mapping, i.e., there exists a matrix norm such that $\|A^{T}\|<1$. A sufficient and equivalent condition is that the spectral radius satisfies $\rho(A)<1$ (since $\rho(A^{T})=\rho(A)$). In the special case where $A$ is diagonal, i.e., $a_{1,2}=a_{2,1}=0$, the mapping $T$ is always a contraction, since
$\|A^{T}\tilde{s}\|\leq a^{*}\|\tilde{s}\|,\qquad a^{*}:=\max\{a_{i,i},\,i=1,\ldots,N\}<1.$
@@
where from hereon we denote $A_{i,j}:=f(a_{1}^{i}\lambda_{1},a_{2}^{j}\lambda_{2})$, $i,j=1,2,\ldots$.
Then, iterating (105) n1n-1 times results in